# Data-Efficient Approach to Humanoid Control via Fine-Tuning a Pre-Trained GPT on Action Data

### v1

@@ Abstract
There are several challenges in developing a model for multi-tasking humanoid control. Reinforcement learning and imitation learning approaches are quite popular in this domain. However, there is a trade-off between the two. Reinforcement learning is not the best option for training a humanoid to perform multiple behaviors due to training time and model size, and imitation learning using kinematics data alone is not appropriate to realize the actual physics of the motion. Training models to perform multiple complex tasks take long training time due to high DoF and complexities of the movements. Although training models offline would be beneficial, another issue is the size of the dataset, usually being quite large to encapsulate multiple movements. Many papers have implemented state of the art deep learning models such as transformers to control humanoid characters and predict their motion based on a large dataset of recorded/reference motion. In this paper, we train a GPT on a large dataset of noisy expert policy rollout observations from a humanoid motion dataset as a pre-trained model and fine tune that model on a smaller dataset of noisy expert policy rollout observations and actions to autoregressively generate physically plausible motion trajectories. We show that it is possible to train a GPT-based foundation model on a smaller dataset in shorter training time to control a humanoid in a realistic physics environment to perform human-like movements.
@@ II-B Transformer based models for humanoid/robot control
Transformers have been shown to be powerful to generate human motion[8, 9]. [10] uses the idea of dual attention to capture spatial and temporal dependencies of known data without relying on hidden states in RNNs or temporal encodings like Discrete Cosine Transformation. This model effectively generates poses that are temporally coherent. MotionGPT fuses language and motion to enhance performance of motion-related tasks such as text-to-motion, motion generation, and motion in-betweening [11, 12]. [13] also demonstrates GPT’s capability to generate motion in a physics engine after taking in one second long motion prompts. In this paper, we train a GPT policy to control a humanoid, similar to [13], but by employing a more data-efficient training strategy.
@@ IV-A Dataset Sizes
Our proposed model was fine-tuned on datasets of half and quarter the size of the Small dataset. We can see that the average generated episode lengths of the HMG trained on quarter and half the size of the Small dataset are lower than that of MoCapAct-Large, which is trained on the Large dataset that is 10 times the size of the Small dataset, and higher than that of MoCapAct-Small. From Fig. 4, we can see that HMG quickly generates higher average prediction lengths on the validation datasets after being trained on the Small dataset, compared to MoCapAct-Large which requires the Large dataset to generate an average prediction length of 5.75 seconds. This shows that our proposed model is significantly data-efficient in learning motion prediction downstream task.
@@ IV-B1 Quantitative Analysis
Using the differences in episode lengths of motion predictions, we were able to find the exact behaviors generated by the models that outperformed the other models in terms of episode lengths. We define an episode length as the total length of an episode for the humanoid to complete the task before episode termination (when it falls down) or the maximum episode length (set to about 15 seconds). We considered the minimum difference in episode lengths of generated behaviors between models to be 6 seconds which we considered significant.
@@ IV-B2 Qualitative Analysis
Avg. Episode
No. of generated motions Length Differences
Model that outlasted the other model (sec)
HMG 9 8.63
MoCapAct-Large 5 8.082
@@ IV-D Motion Prediction Metrics
FID is one of the most important metrics to determine how well the generated motion quality is. The idea behind FID is to compare the distributions between the features from the motion feature extractor based on the real and generated trajectories. The common practice is to train another RNN as a motion feature extractor on the relevant data and then extract the features. Training another RNN in this case is redundant, since the GPT is trained on reconstructing and predicting the motion given the motion prompt, plus, the representation of the features taken from the middle of the GPT network is rich[32]. Hence, HMG and MoCapAct-Large themselves were used as motion feature extractors to calculate FID and DIV respectively. Equation (2) shows how to calculate FID
@@
where $\hat{\mathbf{j}}_{i,T}$ and $\mathbf{j}_{i,T}$ are the generated and real joint poses respectively at the final time step of the trajectory $i$.
@@
Another important metric is the generation diversity (DIV). This metric shows how divergent the generated motion feature distribution are from the real feature motion data. The higher the diversity, the higher the ability of the model to produce novel behaviors. Although, it may be desirable if our model stays true to the real data as much as possible to reduce the risk of generating unpredictable novel behaviors. Equation 5 calculates the diversity score of a model
$\text{DIV}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{E_{d}}(x_{i},\hat{x}_{i})$ (5)
where $x,\hat{x}$ are real and generated features respectively, $n$ is the total number of samples, and $\mathbf{E_{d}}$ is the Euclidean distance.
For our model we desire all the metrics except DIV to be lower than that of MoCapAct-Large. We require ADE and FDE to be low so that the motion similarity is high. We require FID to be low to ensure that the motion quality is as close to that of ground truth. Generally, diversity is preferred to be slightly higher for a multi-tasking model, capable of generating novel and emergent behaviors, but not too much that it deviates largely from the actual motion. From Table IV, we can see that the model displays the desired scores except for DIV. This is probably due to the trade-off of having a low FID score, staying faithful to the ground truth enables the model to generate lesser diverse behaviors, hence the lower DIV score.
TABLE IV: Motion Prediction Scores: $\downarrow$ indicates that lower score is desired and $\uparrow$ indicates that higher score is desired.
Model FID $\downarrow$ ADE $\downarrow$ FDE $\downarrow$ DIV $\uparrow$
HMG 14.012 2.145 0.061 4.291
MoCapAct-Large 20.785 2.443 2.544 4.662
@@ V Discussion
In [13], box plots and histograms were used to evaluate their model’s capability to generate motion. However, these graphs alone are not enough to evaluate the model since it is hard to tell how well the humanoid can actually imitate and complete the ground truth motion prompts accurately. Hence the standard metrics FID, DIV, ADE, and FDE were chosen to additionally evaluate the motion quality, imitation and diversity. These are the standards that have been adopted in [11, 33, 34]. FID, ADE, FDE can evaluate the quality and accuracy of imitation, and DIV is a metric used to evaluate the multi-tasking capability of the model to generate novel behaviors (diversity). The representations of the GPT policies are different when the GPT is trained from scratch and when fine-tuned over a foundation model. Fine-tuning a foundation model gets the model to understand the relationship between the observations themselves and the actions too, therefore learning a much richer representation. This can explain the differences in predicted trajectories observed in Figs. 7-7, and in turn the generation lengths and the other evaluations.
We are curious to see if our foundation model can be trained and fine-tuned on simulation data obtained from other robots such as [35]. If successful, then our foundation model can be fine-tuned to any data that is robot/domain specific. Perhaps we can use this foundation model for other downstream tasks such as motion clustering and action recognition.
Our proposed model is an end-to-end model, therefore separate controllers are not trained to target imitation and balance individually, thus that eliminates the need for a high-level controller for imitation and balance.
@@ VI Conclusion
We proposed a data-efficient approach for motion prediction by pre-training a humanoid motion foundation model on observation data and fine-tuning it on both observation and action data on a motion prediction downstream task. We have evaluated our proposed method’s performance with training from scratch based on several facets; motion prediction metrics, generation lengths, and empirical observations on various cyclic, non-cyclic, and fast paced behaviors. This comprehensive evaluation supports the fact that our model is capable of generating physically plausible trajectories and can perform as well as and even better than the conventional approach. It would be interesting to see how this model can be applied to real humanoid robots by fine-tuning the foundation model on data obtained from various real robots. Furthermore, it would be interesting to see whether the humanoid character’s data from MoCapAct dataset would suffice or the foundation model would need to be trained on real robot data.

### v2

@@ Abstract
There are several challenges in developing a model for multi-tasking humanoid control. Reinforcement learning and imitation learning approaches are quite popular in this domain. However, there is a trade-off between the two. Reinforcement learning is not the best option for training a humanoid to perform multiple behaviors due to training time and model size, and imitation learning using kinematics data alone is not appropriate to realize the actual physics of the motion. Training models to perform multiple complex tasks take long training time due to high DoF and complexities of the movements. Although training models offline would be beneficial, another issue is the size of the dataset, usually being quite large to encapsulate multiple movements. There are few implementations of transformer-based models to control humanoid characters and predict their motion based on a large dataset of recorded/reference motion. In this paper, we train a GPT on a large dataset of noisy expert policy rollout observations from a humanoid motion dataset as a pre-trained model and fine tune that model on a smaller dataset of noisy expert policy rollout observations and actions to autoregressively generate physically plausible motion trajectories. We show that it is possible to train a GPT-based foundation model on a smaller dataset in shorter training time to control a humanoid in a realistic physics environment to perform human-like movements.
@@ II-B Transformer based models for humanoid/robot control
Transformers have been shown to be powerful to generate human motion[8, 9]. [10] uses the idea of dual attention to capture spatial and temporal dependencies of known data without relying on hidden states in RNNs or temporal encodings like Discrete Cosine Transformation. This model effectively generates poses that are temporally coherent. MotionGPT fuses language and motion to enhance performance of motion-related tasks such as text-to-motion, motion generation, and motion in-betweening [11]. [12] also demonstrates GPT’s capability to generate motion in a physics engine after taking in one second long motion prompts. In this paper, we train a GPT policy to control a humanoid, similar to [12], but by employing a more data-efficient training strategy.
@@ IV-A Dataset Sizes
Our proposed model was fine-tuned on datasets of half and quarter the size of the Small dataset. The performances were compared based on episode lengths. We define an episode length as the total length of an episode for the humanoid to complete the task before episode termination (when it falls down) or the maximum episode length (set to about 15 seconds). We can see that the average generated episode lengths of the HMG trained on quarter and half the size of the Small dataset are lower than that of MoCapAct-Large, which is trained on the Large dataset that is 10 times the size of the Small dataset, and higher than that of MoCapAct-Small. From Fig. 4, we can see that HMG quickly generates higher average prediction lengths on the validation datasets after being trained on the Small dataset, compared to MoCapAct-Large which requires the Large dataset to generate an average prediction length of 5.75 seconds. This shows that our proposed model is significantly data-efficient in learning motion prediction downstream task.
@@ IV-B1 Quantitative Analysis
Using the differences in episode lengths of motion predictions, we were able to find the exact behaviors generated by the models that outperformed the other models in terms of episode lengths. We considered the minimum difference in episode lengths of generated behaviors from the validation dataset (which consists of 63 behaviors), between models to be 6 seconds which we considered significant.
@@ IV-B2 Qualitative Analysis
No. of generated motions that Avg. Episode
outlasted the other model Length Differences
Model (from validation dataset) (sec)
HMG 9 out of 63 8.630
MoCapAct-Large 5 out of 63 8.082
@@ IV-D Motion Prediction Metrics
FID is one of the most important metrics to determine how well the generated motion quality is. The idea behind FID is to compare the distributions between the features from the motion feature extractor based on the real and generated trajectories. The common practice is to train another RNN as a motion feature extractor on the relevant data and then extract the features. Training another RNN in this case is redundant, since the GPT is trained on reconstructing and predicting the motion given the motion prompt, plus, the representation of the features taken from the middle of the GPT network is rich[31]. Hence, the pre-trained model which was trained on observation data was used as the motion feature extractor to calculate FID and DIV. Equation (2) shows how to calculate FID
@@
where $\hat{\mathbf{j}}_{i,T}$ and $\mathbf{j}_{i,T}$ are the generated and real joint poses respectively at the final time step of the trajectory $i$. ADE and FDE scores are in radians.
@@
Another important metric is the generation diversity (DIV). This metric shows the variance in the generated motions across all behaviors the model is trained on. It is preferred if the variance is close to that of the real dataset. DIV is calculated by creating two sets of $n$ randomly sampled motions from each generated data and calculating the diversity between the two. Equation 5 calculates the diversity score of a model
$\text{DIV}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{E_{d}}(\mathbf{v_{i}},\mathbf{v_{i}^{\prime}})$ (5)
where $\mathbf{v},\mathbf{v^{\prime}}$ are feature vectors from two sampled sets from the same generated models (expert policy, HMG, MoCapAct-Large) respectively across all behaviors, $n$ is the number of samples in the set, and $\mathbf{E_{d}}$ is the Euclidean distance. We set $n$ to be $200$. We require ADE and FDE to be low so that the motion similarity is high and we require FID to be low to ensure that the motion quality is as close to that of ground truth. From Table IV, we can see that the model displays the desired scores except for DIV which deviates from the real data.
TABLE IV: Motion Prediction Scores: $\downarrow$ indicates that lower score is desired and $\rightarrow$ indicates that score closer to Real (ground truth data) is desired.
Model FID $\downarrow$ ADE $\downarrow$ FDE $\downarrow$ DIV $\rightarrow$
Real 0.000 N/A N/A 7.530
HMG 7.741 6.824 6.616 8.222
MoCapAct-Large 8.415 6.866 6.910 8.151
@@ V Discussion
In [12], box plots and histograms were used to evaluate their model’s capability to generate motion. However, since it is hard to tell how well the humanoid can actually imitate and complete the ground truth motion prompts accurately, the standard metrics FID, DIV, ADE, and FDE were chosen to evaluate the motion quality, imitation and diversity[11, 32, 33]. From our evaluations, we could confirm that our proposed method of fine-tuning a pre-trained model is more data-efficient than training from scratch, and the imitation performance between both the methods are quite similar with our model slightly outperforming the other. However, we cannot control the generation of the motion by conditioning, for instance, if the model generates a motion given a prompt to stand, it may try to either continue to stand or walk. Another limitation is that it uses fully observable state space, whereas realistically we would prefer to only use partially observable variables. Other limitations include imperfect imitation and can probably perform better when fine tuned with reward feedback. We are curious to see if our foundation model can be trained and fine-tuned on simulation data obtained from other robots such as [34].
@@ VI Conclusion
We proposed a data-efficient approach for motion prediction by pre-training a humanoid motion foundation model on observation data and fine-tuning it on both observation and action data on a motion prediction downstream task. We have evaluated our proposed method’s performance with training from scratch based on several facets; training efficiency, motion prediction metrics, generation lengths, and empirical observations on various cyclic, non-cyclic, and fast paced behaviors. We conclude that our proposed method is more data-efficient than the conventional approach.